[Brahms-l] Shift report 20060615 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Thu Jun 15 2006 - 00:00:43 EDT
Shifter leader was PS
Shifters were TE


1. there was one store during the shift but
because no polarization measurement in blue ring
the store is useless for spin physics.
2. took 3 regular data runs: 16132, 16133, 16134, and small one 16135,
after the dumptI took non-TPC pedestal run and tpc pedestal run.
Installed pedestals in the DAQ.
3.No new detector problems encountered. 
_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Jun 15 00:01:15 2006

This archive was generated by hypermail 2.1.8 : Thu Jun 15 2006 - 00:01:23 EDT